OpenStudy (anonymous):

integrate sinh^2(x) cosh^3 (x) dx

OpenStudy (anonymous):

First thing's first: Hyperbolic Trig Identities (YAY!) Recall:\[\cosh ^{2}(x)-\sinh ^{2}(x)=1\] Which form of this identity should we use to integrate this problem? Hint: Think of the u substitution you want to make later on.

OpenStudy (anonymous):

so it will be like this \[\int\limits \sinh^2 x \cosh x(\sinh^2 x+ 1)\] then let u=sinhx

OpenStudy (anonymous):

Exactly. Continue, you're doing great!

OpenStudy (anonymous):

\[du=\cosh x\]\[u^2 coshx(u^2 +1) \frac{ du }{ coshx }\] \[\int\limits u^2(u^2+1)\]

OpenStudy (anonymous):

Don't forget your du at the end, and then you can just integrate and replace your u and you have your answer.

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

I am glad to help, but did the entire problem. Nice work!

OpenStudy (anonymous):

*but you did the entire problem.

OpenStudy (anonymous):

i got the answer actually but since the answer is different from answer sheet, so i want confirmation.

OpenStudy (solomonzelman):

You are my hyperbolic expert! :o

OpenStudy (anonymous):

Well next time you can just post problem and your answer and someone will double check it if you want to skip me (or someone else) trying to explain all of the steps :P

OpenStudy (anonymous):

noted, thanks again

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \frac{u^5}{5} +\frac{u^3}{3}+C}\) Recall that you did \(\color{#000000 }{ \displaystyle u=\sinh(x)}\) So, to sub back Do you get a different answer?

OpenStudy (solomonzelman):

What answer did you get?

OpenStudy (anonymous):

yeah i got the same answer, but the sheet answer is \[-\frac{ sinhx }{ 8 } + \frac{ 1 }{ 4 }\sinh3x +\frac{ 1 }{ 80 }\sinh5x\]

OpenStudy (anonymous):

It doesn't even have a C :( The horror!

OpenStudy (anonymous):

lol, i know right

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \frac{\sinh^5(x) }{5}+\frac{\sinh^3(x) }{3} +C }\) that is what I seem to get - don't yet know if I can conviniently simplify this, if I could simplify this at all...

OpenStudy (anonymous):

yeah that's the answer i got, dont know how on earth the sheet answer its so complex

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 1-\sinh^2(x)\ne \cosh^2(x) ~~\forall x}\)

OpenStudy (anonymous):

oh yeah i mislook that. thank you

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \sinh^2(x)+\cosh^2(x)~~~~\square~~~~1}\) \(\color{#000000 }{ \displaystyle \frac{1}{4}(e^x-e^{-x})^2+\frac{1}{4}(e^x+e^{-x})^2 ~~~~\square~~~~1}\) \(\color{#000000 }{ \displaystyle (e^x-e^{-x})^2+(e^x+e^{-x})^2 ~~~~\square~~~~4}\) \(\color{#000000 }{ \displaystyle (e^{2x}-2+e^{-2x})+(e^{2x}+2+e^{-2x}) ~~~~\square~~~~4}\) \(\color{#000000 }{ \displaystyle 2e^{2x}+2e^{-2x} ~~~~\square~~~~4}\)

OpenStudy (solomonzelman):

Again, using the definitions are very helpful, and the lesson is that hyperbolic functions don't (always) abide the same rules are regular trig functions do.

OpenStudy (solomonzelman):

I don't know any hyperbolic properties; I haven't ever integrated them really. But, if I apply defs for sinh(x) and coh(x) then I can get some answer, although it might be painful and will definitely differ

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \int \sinh^2(x) \cosh^3 (x) dx }\) \(\color{#000000 }{ \displaystyle 2^5\int \left(e^{2x}+e^{-2x}-2\right)\left(e^{2x}+e^{-2x}+2\right)\left(e^{x}+e^{-x}\right) dx }\) \(\color{#000000 }{ \displaystyle 32\int \left(\frac{e^{4x}+2e^{2x}-1}{e^{2x}}\right)\left(\frac{e^{4x}+2e^{2x}+1}{e^{2x}}\right)\left(\frac{e^{2x}+1}{e^x} \right) dx }\) \(\color{#000000 }{ \displaystyle 32\int \frac{1}{e^{5x}}\left(e^{4x}+2e^{2x}-1\right)\left(e^{4x}+2e^{2x}+1\right)\left(e^{2x}+1 \right) dx }\) I honestly would do this if I were on the test an didn't have the web to look for rules....

OpenStudy (anonymous):

in our test the identities will be given, so one less burden to us, but thanks for the alternative way